# SLO-aware GPU Frequency Scaling for Energy Efficient LLM Inference Serving

### v1

@@ V-C KV cache and Batch size Projection Mechanism Evaluation
VI Related Work
@@
LLM optimization is an active research topic, with many works aiming to improve the inference process [74].
@@
LLM inference optimization: Recent research has explored various optimization techniques for LLM inference. These efforts address different facets of LLM inference, including i) efficient request batching [19, 65] with inflight batching introduced by Orca [65] being a pioneering work in the field, ii) efficient LLM serving and request scheduling [65, 32, 4, 50, 35, 61, 21, 43, 72, 52, 40, 58] with NVIDIA’s Triton [40] being a state-of-the-art solution for inference serving offering optimized performance, iii) LLM memory optimization [51, 28, 19, 70] with paged attention [28] being a notable innovation in the area, iv) model scaling and parallelism on multiple GPUs [6, 45, 23], v) speculative decoding and sampling [63, 34, 53, 9, 30, 62], vi) parameter sharing [73] and vii) optimization of GPU kernels [57, 66, 16, 15, 2, 6] with Flash Attention [16, 15] and NVIDIA’s FasterTransformer  [2] being the most notable works. While these works primarily target performance optimization, neglecting energy efficiency, they offer orthogonal techniques that can complement throttLL’eM. Notably, as mentioned in Sec. V-A, throttLL’eM is built on top of Triton [40] which integrates optimizations like inflight batching [65], paged attention [28] and flash attention [15].
@@
LLM implications on Energy/Power: Recent works explore the energy and power implications of LLM inference. Studies like [48, 54, 11] characterize energy consumption and carbon emissions but lack frameworks for energy-efficient LLM inference. The authors of [60] model the performance and energy consumption of LLM queries using an analytical approach. They use this model to solve an offline scheduling optimization problem. However, they overlook the dynamic nature of inference systems and KV cache implications (cf. Sec. III-B). In [25] a framework for energy-efficient LLM inference is proposed, but is limited to fixed-length queries (max_tokens) on single-GPU models. Lastly, [41] presents POLCA, a framework for power oversubscription in LLM clusters. throttLL’eM can complement POLCA by performing frequency scaling under power-capping constraints, enhancing the margins for power oversubscription.
VII Conclusion
@@
LLM inference presents an emerging workload with considerable energy footprint. In this work we performed an extensive study on the factors that affect the inference process, gathering insights on performance and energy efficiency. Guided by these insights, we designed throttLL’eM, a framework that uses a predictive, ML-driven approach to increase LLM energy efficiency under SLOs. throttLL’eM improves efficiency by over 1.71×1.71$\times$1.71 ×, consuming up to $43.8$% less energy than Triton.

### v2

@@ 5.3 KV cache and Batch size Projection Mechanism Evaluation
Figure 12: a) Reduction in iterations and b) Increase in avg. batch size of throttLL’eM to autoscale-only for all error levels.
Figure 13: Comparison of a) $p99^{th}$ E2E and b) avg. Power for different implementations at increasing system load levels.
5.6 Extended Evaluation & Comparison
5.6.1 Sources of Energy Gains
The observed energy gains in throttLL’eM are driven by two key mechanisms. First, autoscaling adjusts the allocated capacity, resulting in a 20.8% reduction in energy consumption compared to Triton TP4, as fewer GPUs are required, lowering static power draw across the entire trace. Second, frequency scaling further reduces energy consumption by operating GPUs at lower frequencies, as shown in Sec. 3.1.1. A side effect of this is that lower frequencies slow down the inference process, allowing more requests to accumulate in each batch, thereby artificially increasing batch size. This improves energy efficiency (cf. Fig. 2e), but also decreases the number of performed LLM iterations, leveraging the higher batch-level parallelism to generate more tokens per forward pass. Figure 12 illustrates this behaviour, showing how throttLL’eM increases batch size and reduces iterations compared to autoscaling alone. At 0% predictor error level, throttLL’eM manages to increase batch size by a geo-mean of $24.5\%$ across all engines, performing $24.3\%$ less iterations. Although higher errors result in higher frequencies, which diminish these benefits, the overall trend remains consistent.
5.6.2 throttLL’eM operation under varying loads
We examine the trade-off between energy efficiency and latency by progressively increasing the system load from 10% to 100% across different implementations. Figure 13b demonstrates how throttLL’eM is able to reduce average power compared to Triton and autoscaling-only by downsizing engines and reducing frequencies. In this manner, throttLL’eM trades-off latency, which as shown in Fig. 13a, is brought close to that of peak load ($31.3$s per Tab. 2). While autoscaling-only can significantly reduce the average power draw, e.g., at 10% load level, in other cases the frequency scaling logic of throttLL’eM offers additional gains which cannot be achieved with coarse-grained optimizations like autoscaling.
Figure 14: a) $p99^{th}$ E2E and b) Avg. power draw of throttLL’eM (w/o autoscaling) and a Retail-inspired implementation.
5.6.3 Comparison with other DVFS schemes
To demonstrate the effectiveness of our LLM-aware design choices, we compare throttLL’eM with a SotA microservice-based DVFS scheme, similar to ReTail [retail]. ReTail automatically identifies input and application features that strongly correlate to E2E latency, in our case: frequency, batch size, KV cache utilization and input/output length. Using these, it trains a linear regression model to predict latency and selects the lowest frequency that satisfies the SLOs for all requests in the queue. Upon receiving a new request, ReTail evaluates available frequencies and chooses the minimum that meets SLO requirements.
To ensure a fair comparison, we used the LLaMa2-13B model in its TP4 configuration, as ReTail does not support autoscaling. Both implementations were tested at a 30% predictor error level. Figure 14b shows that throttLL’eM consistently achieves lower power consumption, while Fig. 14a demonstrates its ability to more aggressively approach the latency deadlines. We attribute this to i) throttLL’eM’s capability to forecast KV cache utilization and batch size, and ii) a more accurate ML-based performance prediction model. In detail, ReTail relies on the currently known values of KV cache utilization and batch size for latency prediction, which however change over time, whereas throttLL’eM uses forecasted values for future iterations. Moreover, our analysis showed that ReTail’s linear regression model struggles with E2E latency prediction, achieving $R^{2}$ close to 0.4. This is due to the non-linear E2E latency slowdown, as token generation proceeds. throttLL’eM addresses this by using iteration-level predictions which are then aggregated to produce accurate E2E latency predictions. Overall, across the entire trace throttLL’eM consumes 11.911.9% lower energy, a value that would have been higher had autoscaling also been enabled.
5.6.4 Evaluation on different systems/LLMs
We further examine how throttLL’eM adapts to diverse systems and LLMs, by deploying a Mixtral 8x7B Mixture-of-Experts LLM [jiang2024mixtral] on a g5.48xlarge AWS instance with $8\times$ NVIDIA A10G 24GB GPUs. This system features lower performance than the previously examined one, lacking the HBM2 memory and the 600 GB/s NVLink interconnects of p4d.24xlarge. After profiling and training, $\mathcal{M}$ features an $R^{2}$ (MAPE) score of $0.98$ ($5.8\%$), displaying accurate modelling performance. We scale the trace at the peak load of 1.25 RPS and a target latency of 30s. Figure 15 displays the $p99^{th}$ E2E latencies and the average power draw of both implementations when serving the scaled trace. In this setup, throttLL’eM is able to reduce average power by $35.9\%$, while still featuring a $p99^{th}$ E2E latency of $28.4$s, being well within the $30$s SLO for this configuration.
Figure 15: a) E2E latency b) Total Energy Consumption c) Total iterations and d) Average Batch size of Triton (left) and throttLL’eM (right) of Mixtral 8x7B on NVIDIA A10G (TP8).
6 Related Work
@@
LLM serving optimization is an active research topic, with many works aiming to improve the inference process [zhou2024survey].
@@
LLM inference optimization: Recent research has explored various optimization techniques for LLM inference. These efforts address different facets of LLM inference, including i) efficient request batching [fang2021turbotransformers, gyeong2022orca] with inflight batching introduced by Orca [gyeong2022orca] being a pioneering work in the field, ii) efficient LLM serving and request scheduling [gyeong2022orca, li2023alpaserve, agrawal2023sarathi, sheng2024fairness, spotserve, wu2023fast, hu2024inference, patel2024splitwise, zhong2024distserve, song2023powerinferfastlargelanguage, triton-inference-server, wang2023tabi] with NVIDIA’s Triton [triton-inference-server] being a state-of-the-art solution for inference serving offering optimized performance, iii) LLM memory optimization [sheng2023flexgen, kwon2023pagedattention, fang2021turbotransformers, zhang2024h2o] with paged attention [kwon2023pagedattention] being a notable innovation in the area, iv) model scaling and parallelism on multiple GPUs [aminabadi2022deepspeed, pope2023efficiently, jiand2024megascale], v) speculative decoding and sampling [xia2024unlocking, miao2024specinfer, stern2018blockwise, chen2023accelerating, leviathan2023fast, xia-etal-2023-speculative], vi) parameter sharing [zhou2022pets] and vii) optimization of GPU kernels [wang-etal-2021-lightseq, zhai2023bytetransformer, dao2022flashattention, dao2023flashattention, faster_transformer, aminabadi2022deepspeed] with Flash Attention [dao2022flashattention, dao2023flashattention] and NVIDIA’s FasterTransformer [faster_transformer] being the most notable works. The work presented in [SoCC-LLM-queues] manages multiple queues for interactive and batched inference requests, avoiding potential head-of-line blocking. While these works primarily target performance optimization, neglecting energy efficiency, they offer orthogonal techniques that can complement throttLL’eM. Notably, as mentioned in Sec. 5.1, throttLL’eM is built on top of Triton [triton-inference-server] which integrates optimizations like inflight batching [gyeong2022orca], paged attention [kwon2023pagedattention] and flash attention [dao2023flashattention].
@@
Power and Energy Efficient LLM inference serving: Recent works explore the energy and power implications of LLM inference. Studies such as  [samsi2023words, stojkovic2024towards, chien2023reducing] characterize energy consumption and carbon emissions but lack frameworks for energy-efficient LLM inference. The authors of [wilkins2024offline] model the performance and energy consumption of LLM queries using an analytical approach. They use this model to solve an offline scheduling optimization problem. However, they overlook the dynamic nature of inference systems and KV cache implications (cf. Sec. 3.2). In [kakolyris-CAL] a framework for energy-efficient LLM inference is proposed, but is limited to fixed-length queries (max_tokens) on single-GPU models. A concurrent work, µ-Serve [u-serve-atc] performs intelligent model partition placement and introduces a multiplicative-increase-additive-decrease (MIAD)-based DVFS scheme to reduce power consumption while meeting the SLOs of LLM requests based on their predicted generation length and SLO slack. Lastly, [patel2023polca] presents POLCA, a framework for power oversubscription in LLM clusters. throttLL’eM can complement POLCA by performing frequency scaling under power-capping constraints, enhancing the margins for power oversubscription.
7 Conclusion
@@
LLM inference presents an emerging workload with considerable energy footprint. In this work we performed an extensive study on the factors that affect the inference process, gathering insights on performance and energy efficiency. Guided by these insights, we designed throttLL’eM, a framework that uses a predictive, ML-driven approach to increase the energy efficiency of LLM inference serving under SLOs. throttLL’eM improves efficiency by over $1.71\times$, consuming up to 43.843.8% less energy than Triton.
Acknowledgements
The authors would like to thank the National Infrastructure for Research and Technology Network (GRNET) for funding the cloud infrastructure used to evaluate throttLL’eM. They further acknowledge Georgios Zervakis and Panagiotis Hadjidoukas for granting access to the DGX Deep Learning System of the Computer Engineering and Informatics Department (CEID), University of Patras, utilized during development. Lastly, they are grateful to Konstantinos Kanellopoulos and the anonymous reviewers for their constructive feedback, which enhanced the quality of this paper.
This work has been partially funded by the PRIVATEER project. PRIVATEER has received funding from the Smart Networks and Services Joint Undertaking (SNS JU) under the European Union’s Horizon Europe research and innovation programme under Grant Agreement No. 101096110. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or SNS JU. Neither the European Union nor the granting authority can be held responsible for them.